OpenStudy (anonymous):

write (6x^2-13x-5)/(9x^3-x) in lowest terms

OpenStudy (lgbasallote):

\[\Large \frac{6x^2 - 14x - 5}{9x^3 - x} \implies \frac{(3x+1)(2x-5)}{x(9x^2 - 1)} \implies \frac{(3x+1)(2x-5)}{x(3x +1)(3x-1)}\] does that help?

OpenStudy (anonymous):

not reLLY cant see the whole thing

OpenStudy (anonymous):

its 6x^2-13x-5

OpenStudy (lgbasallote):

uhmm yeah..that was just a typo hehe

OpenStudy (lgbasallote):

the next steps followed the 13x...i jsut wrote 13x wrongly

OpenStudy (lgbasallote):

so do you get it? i just factored out the numerator and denominator

OpenStudy (anonymous):

well the last part i cant see for some reason hehe

OpenStudy (anonymous):

what happen after is factored?

OpenStudy (lgbasallote):

the last part is \[\Large \frac{(3x+1)(2x-5)}{x(3x+1)(3x-1)}\]

OpenStudy (lgbasallote):

do you see what you're going to do now?

OpenStudy (anonymous):

elimnate right

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

then you have the answer

OpenStudy (anonymous):

gracias

OpenStudy (lgbasallote):

de nada